A car accelerates from rest at 2 m/s/s. After 10 seconds the car has traveled?
Answer #1

initial velocity =u= 0m/sec

accelaration = a = 2m/s2

using the second equation of motion

S=ut+1/2at2

u is given to be zero

therefore s=1/2at2 = (1/2) *2*100 =100 metres